OpenStudy (anonymous):

]Choose the equation below that represents the line passing through the point (2, -5) with a slope of -3. y = -3x - 13 y = -3x + 11 y = -3x + 13 y = -3x + 1

OpenStudy (anonymous):

use y=mx+b where m is the slope start plugging in values to find b -5=-3*2+b and b=1 so now you have all the ingredients

OpenStudy (anonymous):

y=-3x+1 so answer is D